OpenStudy (anonymous):

find the values of x for which the equation cos x=-1 is true. let k represent the integer

OpenStudy (anonymous):

cos x=-1 x = cos^-1 (-1) = 180* for [0,360] cycle x = 180* + 360k where k represents an integer.

OpenStudy (anonymous):

the only options they give me are a.) 2pi(k) b.)pi/2+2pi(k) c.)pi+2pi(k) d.) 3pi/2+2pi(k)

OpenStudy (anonymous):

cosine is -1 at \(\pi,3\pi,5\pi,...\)

OpenStudy (anonymous):

odd numbers in other words, times \(\pi\)

OpenStudy (anonymous):

go with choice C (it is always C)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

each roller under a conveyor belt has a radius of 0.5 meters. the rollers turn at a rate of 30 revolutions per minute. what is the linear velocity of the conveyor belt